# Using Dynamic Safety Margins as Control Barrier Functions

### v2

@@ 2.2 Control Barrier Functions
$\beta(\boldsymbol{\mathbf{x}})=\underset{\boldsymbol{\mathbf{u}}\in\mathcal{K}(\boldsymbol{\mathbf{x}})}{\operatorname{argmin}}\>\|\boldsymbol{\mathbf{u}}-\kappa(\boldsymbol{\mathbf{x}})\|^{2},$ (12)
and $\kappa:\mathbb{X}\to\mathbb{U}$ is a nominal control policy with desirable closed-loop performance. When (9) holds, the optimization problem (12) is always feasible, making $\beta$ well-defined. If, in addition, the policy $\beta$ is locally Lipschitz, we can guarantee the existence and uniqueness of a solution $\boldsymbol{\mathbf{x}}(t)$ to the closed-loop system (11) and that it will satisfy $\boldsymbol{\mathbf{x}}(t)\in\mathcal{X}$. In upcoming sections, we will specify conditions under which $\beta$ is locally Lipschitz.
As noted in the introduction, finding a CBF for arbitrary constraint sets $\mathcal{X}\subset\mathbb{X}$ is a challenging task. In fact, Lemma 1 states that, when $\mathcal{U}$ is compact, a CBF $h$ with a well-defined policy $\beta$ exists only if $\mathcal{X}$ is control invariant, which is generally not the case. In spite of this, many practitioners insist on imposing $\mathcal{C}=\mathcal{X}$ to construct a “candidate” CBF (i.e. a continuously differentiable function $h:\mathcal{D}\to\mathbb{R}$ that satisfies $\mathcal{X}=\{\boldsymbol{\mathbf{x}}\in\mathcal{D}:h(\boldsymbol{\mathbf{x}})\geq 0\}$) and then tuning a class $\mathscr{K}_{\infty}$ function $\alpha$ on a case-by-case basis. Although this approach can yield good performance, it suffers from the possibility that $\mathcal{K}(\boldsymbol{\mathbf{x}})=\emptyset$ for some $x\in\mathcal{X}$, which makes the controller $\beta$ ill-defined.
In this paper, we leverage results from the ERG framework to design a valid CBF that satisfies Definition 2 and, additionally, is guaranteed to satisfy (9).
@@ 2.3 Dynamic Safety Margins
Dynamic safety margins are real-valued functions that quantify the distance to constraint violation for a prestabilized system subject to a constant reference. In order to use them, we must make the following assumptions about system (1).
@@
System (1) admits a simply connected equilibrium manifold parameterized by continuous functions $\bar{x}:\mathbb{V}\to\mathbb{X}$ and $\bar{u}:\mathbb{V}\to\mathbb{U}$ such that
$\forall\boldsymbol{\mathbf{v}}\in\mathbb{V},\quad f\big{(}\bar{x}(\boldsymbol{\mathbf{v}})\big{)}+g\big{(}\bar{x}(\boldsymbol{\mathbf{v}})\big{)}\bar{u}(\boldsymbol{\mathbf{v}})=0,$ (13)
where $\mathbb{V}\subset\mathbb{R}^{l}$ is the reference space. Moreover, every equilibrium point $\bar{x}(\boldsymbol{\mathbf{v}})$ is stabilizable.
The vector $\boldsymbol{\mathbf{v}}\in\mathbb{V}$ is hereafter referred to as a “reference” for the system because it identifies a specific equilibrium point. The set of steady-state admissible references is
$\mathcal{V}=\{\boldsymbol{\mathbf{v}}\in\mathbb{V}\mid\bar{x}(\boldsymbol{\mathbf{v}})\in\mathcal{X},~{}\bar{u}(\boldsymbol{\mathbf{v}})\in\mathcal{U}\}.$ (14)
Note that, since $\bar{x}$, $\bar{u}$ are continuous and $\mathcal{X}$, $\mathcal{U}$ are closed sets, $\mathcal{V}$ is closed in $\mathbb{V}$.
@@
The set $\mathcal{V}\subset\mathbb{V}$ is not empty, i.e., system (1) features at least one steady-state admissible equilibrium point.
In essence, Assumption 1 states that the system has an open set of stabilizable equilibria, and Assumption 2 states that at least one of the equilibria is safe. By Assumption 1, let $\pi:\mathbb{X}\times\mathbb{V}\to\mathbb{U}$, locally Lipschitz continuous on $\mathbb{X}$ and continuous on $\mathbb{V}$, be a prestabilizing control law such that $\pi\big{(}\bar{x}(\boldsymbol{\mathbf{v}}),\boldsymbol{\mathbf{v}}\big{)}=\bar{u}(\boldsymbol{\mathbf{v}})$, and $\bar{x}(\boldsymbol{\mathbf{v}})$ is an asymptotically stable equilibrium point of the prestabilized system
$\dot{\boldsymbol{\mathbf{x}}}=f(\boldsymbol{\mathbf{x}})+g(\boldsymbol{\mathbf{x}})\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\triangleq f_{\pi}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}).$ (15)
Given a reference $\boldsymbol{\mathbf{v}}\in\mathbb{V}$, the prestabilized dynamics $f_{\pi}$ ensure the existence of an open set $\mathcal{D}_{\boldsymbol{\mathbf{v}}}\subset\mathbb{X}$ that satisfies $\bar{x}(\boldsymbol{\mathbf{v}})\in\mathcal{D}_{\boldsymbol{\mathbf{v}}}$ and is such that
@@
The set $\mathcal{D}_{\boldsymbol{\mathbf{v}}}$ is a region of attraction for $\bar{x}(\boldsymbol{\mathbf{v}})$. We assume small changes in $\boldsymbol{\mathbf{v}}$ yield small changes in $\mathcal{D}_{\boldsymbol{\mathbf{v}}}\subset\mathbb{X}$. Formally,
@@
The set
$\mathcal{D}_{\boldsymbol{\mathbf{v}}}\times\mathbb{V}\triangleq\{(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\mathbb{X}\times\mathbb{V}\mid\boldsymbol{\mathbf{x}}\in\mathcal{D}_{\boldsymbol{\mathbf{v}}}\},$ (17)
is open in $\mathbb{X}\times\mathbb{V}$.
Note that, for each reference $\boldsymbol{\mathbf{v}}\in\mathbb{V}$, the input constraint set $\mathcal{U}$ induces state constraints on the prestabilized system. This allows us to define the reference-dependent state constraint set
By continuity of $\pi$ on $\mathbb{V}$, it follows that $\mathcal{X}_{\boldsymbol{\mathbf{v}}}$ is closed in $\mathbb{X}$. The following DSM definition satisfies the one proposed in [22, Definition 1], although it makes stronger assumptions by requiring a) differentiability rather than continuity and b) invariance rather than strong returnability.
@@
A continuously differentiable function $\Delta:\!\mathcal{D}_{\boldsymbol{\mathbf{v}}}\!\times\mathbb{V}\!\!\to\!\!\mathbb{R}$ is a dynamic safety margin (DSM) for $\pi$ if, $\forall\boldsymbol{\mathbf{v}}\!\in\!\mathbb{V}$,
$\boldsymbol{\mathbf{v}}\in\mathcal{V}~{}$ $\implies~{}\Delta\big{(}\bar{x}(\boldsymbol{\mathbf{v}}),\boldsymbol{\mathbf{v}}\big{)}\geq 0,$ (19a)
$\Delta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq 0~{}$ $\implies~{}\boldsymbol{\mathbf{x}}\in\mathcal{X}_{\boldsymbol{\mathbf{v}}},$ (19b)
$\Delta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq 0~{}$ $\implies~{}\frac{\partial\Delta}{\partial\boldsymbol{\mathbf{x}}}f_{\pi}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq 0,$ (19c)
$\epsilon>0~{}$ $\implies~{}\mathcal{C}_{\boldsymbol{\mathbf{v}}}^{\epsilon}~{}\text{is closed in $\mathbb{X}$,}$ (19d)
where $\mathcal{C}_{\boldsymbol{\mathbf{v}}}^{\epsilon}=\{\boldsymbol{\mathbf{x}}\in\mathcal{D}_{\boldsymbol{\mathbf{v}}}\mid\Delta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq\epsilon\}$.
Lemma 2.5.
If $\Delta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq 0$, then $\boldsymbol{\mathbf{v}}\in\mathcal{V}$.
Proof 2.6.
Let $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\mathcal{D}_{\boldsymbol{\mathbf{v}}}\times\mathbb{V}$ be such that $\Delta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{r}})\geq 0$. Given the initial condition $\boldsymbol{\mathbf{x}}(0)=\boldsymbol{\mathbf{x}}$, let $\dot{\boldsymbol{\mathbf{x}}}(t)=f_{\pi}\big{(}\boldsymbol{\mathbf{x}}(t),\boldsymbol{\mathbf{v}}\big{)}$. Then, it follows from the continuity of $\Delta$ and property (19c) that $\Delta\big{(}\boldsymbol{\mathbf{x}}(t),\boldsymbol{\mathbf{v}}\big{)}\geq 0$. Therefore, by property (19b), $\boldsymbol{\mathbf{x}}(t)\in\mathcal{X}_{\boldsymbol{\mathbf{v}}}$ for all $t\geq 0$. Taking the limit as $t\to\infty$, we obtain
$\lim_{t\to\infty}\boldsymbol{\mathbf{x}}(t)=\bar{x}(\boldsymbol{\mathbf{v}})\in\mathcal{X}_{\boldsymbol{\mathbf{v}}}\subset\mathcal{X}.$ (20)
Finally, $\pi\big{(}\bar{x}(\boldsymbol{\mathbf{v}}),\boldsymbol{\mathbf{v}}\big{)}=\bar{u}(\boldsymbol{\mathbf{v}})\in\mathcal{U}$, and $\boldsymbol{\mathbf{v}}\in\mathcal{V}$.
@@
$\begin{bmatrix}\dot{\boldsymbol{\mathbf{x}}}\\ \dot{\boldsymbol{\mathbf{v}}}\end{bmatrix}=\begin{bmatrix}f(\boldsymbol{\mathbf{x}})+g(\boldsymbol{\mathbf{x}})\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\\ \Delta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\rho(\boldsymbol{\mathbf{v}},\boldsymbol{\mathbf{r}})\end{bmatrix},$ (21)
where $\rho:\mathbb{V}\times\mathbb{V}\to\mathbb{R}^{l}$ is any locally Lipschitz continuous function such that $\boldsymbol{\mathbf{r}}\in\mathbb{V}$ is an asymptotically stable equilibrium point of the system $\dot{\boldsymbol{\mathbf{v}}}(t)=\rho\big{(}\boldsymbol{\mathbf{v}}(t),\boldsymbol{\mathbf{r}}\big{)}$. In the ERG literature, $\rho$ is called a navigation field [22, Definition 2]. System (21) ensures that, if $\big{(}\boldsymbol{\mathbf{x}}(0),\boldsymbol{\mathbf{v}}(0)\big{)}\in\mathcal{X}_{\boldsymbol{\mathbf{v}}(0)}\times\mathcal{V}$, then $\big{(}\boldsymbol{\mathbf{x}}(t),\boldsymbol{\mathbf{v}}(t)\big{)}\in\mathcal{X}_{\boldsymbol{\mathbf{v}}(t)}\times\mathcal{V}$ at all times [22]. The ERG literature provides systematic tools for computing DSMs for various classes of systems [28, 22, 29, 30, 31]. The following section shows how these tools from the ERG literature can be used to systematically construct valid CBFs.

### v3

@@ 2.2 Control Barrier Functions
$\beta(\boldsymbol{\mathbf{x}})=\underset{\boldsymbol{\mathbf{u}}\in\mathcal{K}_{\alpha}(\boldsymbol{\mathbf{x}})}{\operatorname{argmin}}\>\|\boldsymbol{\mathbf{u}}-\kappa(\boldsymbol{\mathbf{x}})\|^{2},$ (12)
and $\kappa:\mathbb{R}^{n}\to\mathbb{R}^{m}$ is a nominal control policy with desirable closed-loop performance. When (9) holds, the optimization problem (12) is always feasible, making $\beta$ well-defined. If, in addition, the policy $\beta$ is locally Lipschitz continuous, there exists a unique solution $\boldsymbol{\mathbf{x}}(t)$ to the closed-loop system (11), and it satisfies $\boldsymbol{\mathbf{x}}(t)\!\in\!\mathcal{C}\!\subset\!\mathcal{X}$. In upcoming sections, we provide conditions under which $\beta$ is locally Lipschitz continuous.
Often, the sets $\mathcal{X}$ and $\mathcal{U}$ are the intersection of multiple state and input constraints. In these cases, it is usually easier to design a separate CBF for each constraint and then enforce all the conditions in one CBF-based program. Unfortunately, there is typically no guarantee that feasibility is retained when all the constraints are superimposed. The following property certifies feasibility is retained in these cases
Definition 2.4 ([9, Definition 2]).
Consider $n_{c}\in\mathbb{N}$ continuously differentiable functions $h_{i}:\mathcal{D}_{i}\to\mathbb{R}$ and let $\mathcal{C}_{i}=\{\boldsymbol{\mathbf{x}}\in\mathcal{D}_{i}\mid h_{i}(\boldsymbol{\mathbf{x}})\geq 0\}\subset\mathbb{R}^{n}$, with $\mathcal{C}=\cap_{i=1}^{n_{c}}\mathcal{C}_{i}$. The functions $h_{i}$ are said to have the control-sharing property if there exists a class $\mathscr{K}_{\infty}$ function $\alpha$ satisfying $\forall\boldsymbol{\mathbf{x}}\in\mathcal{C}$, $\exists\boldsymbol{\mathbf{u}}\in\mathcal{U}$, such that
$\min_{i\in\{1,\ldots,n_{c}\}}\left[L_{f}h_{i}(\boldsymbol{\mathbf{x}})+L_{g}h_{i}(\boldsymbol{\mathbf{x}})\boldsymbol{\mathbf{u}}+\alpha\big{(}h_{i}(\boldsymbol{\mathbf{x}})\big{)}\right]\!\geq\!0.$ (13)
In this paper, we call a smooth function $h:\mathcal{D}\to\mathbb{R}^{n_{c}}$ whose channels $h_{i}$ have the control-sharing property a (vector-valued) control barrier function.
The functions $h_{i}$ in Definition 2.4 are deliberately not labeled as CBFs individually (as opposed to [9, Definition 2]) because such a requirement is unnecessarily limiting. In fact, if 00 is a regular value of each $h_{i}$, Lemma 1 also holds for the set $\mathcal{C}$ defined as the intersection of the $\mathcal{C}_{i}$ even if each $\mathcal{C}_{i}$ is individually not compact (see [28, Proposition 4.3.7] for more details.) The optimization-based policy $\beta$ can also be used with (vector-valued) CBFs with the following admissible input set
[…]
As noted in the introduction, finding a CBF for arbitrary constraint sets $\mathcal{X}\subset\mathbb{R}^{n}$ is a challenging task. In fact, Lemma 1 states that, when $\mathcal{U}$ is compact, a CBF $h:\mathcal{D}\to\mathbb{R}$ satisfying $\mathcal{X}=\{\boldsymbol{\mathbf{x}}\in\mathcal{D}\mid h(\boldsymbol{\mathbf{x}})\geq 0\}$ with a well-defined policy $\beta$ exists only if $\mathcal{X}$ is control invariant, which is generally not the case. In spite of this, many practitioners insist on imposing $\mathcal{C}=\mathcal{X}$ to construct a “candidate” CBF (i.e. a continuously differentiable function $h:\mathcal{D}\to\mathbb{R}$ that satisfies $\mathcal{X}=\{\boldsymbol{\mathbf{x}}\in\mathcal{D}\mid h(\boldsymbol{\mathbf{x}})\geq 0\}$) and then tuning a class $\mathscr{K}_{\infty}$ function $\alpha$ on a case-by-case basis. Although this approach can yield good performance, it suffers from the possibility that $\mathcal{K}_{\alpha}(\boldsymbol{\mathbf{x}})=\emptyset$ for some $\boldsymbol{\mathbf{x}}\in\mathcal{X}$, which makes the controller $\beta$ ill-defined.
In this paper, we leverage results from the ERG framework to design valid CBFs that satisfy Definition 2 (and more generally Definition 2.4) and, additionally, are guaranteed to satisfy (9).
@@ 2.3 Dynamic Safety Margins
Dynamic safety margins are functions that quantify the distance to constraint violation for a prestabilized system subject to a constant reference. In order to use them, we must make the following assumptions about system (1).
@@
System (1) admits a simply connected equilibrium manifold parameterized by continuous functions $\bar{x}:\mathbb{R}^{l}\to\mathbb{R}^{n}$ and $\bar{u}:\mathbb{R}^{l}\to\mathbb{R}^{m}$ such that
$\forall\boldsymbol{\mathbf{v}}\in\mathbb{R}^{l},\quad f\big{(}\bar{x}(\boldsymbol{\mathbf{v}})\big{)}+g\big{(}\bar{x}(\boldsymbol{\mathbf{v}})\big{)}\bar{u}(\boldsymbol{\mathbf{v}})=0.$ (15)
Moreover, every equilibrium point $\bar{x}(\boldsymbol{\mathbf{v}})$ is stabilizable.
The vector $\boldsymbol{\mathbf{v}}\in\mathbb{R}^{l}$ is hereafter referred to as a “reference” for the system because it identifies a specific equilibrium point. The set of steady-state admissible references is
$\mathcal{V}=\{\boldsymbol{\mathbf{v}}\in\mathbb{R}^{l}\mid\bar{x}(\boldsymbol{\mathbf{v}})\in\mathcal{X},~{}\bar{u}(\boldsymbol{\mathbf{v}})\in\mathcal{U}\}.$ (16)
Note that, since $\bar{x}$, $\bar{u}$ are continuous and $\mathcal{X}$, $\mathcal{U}$ are closed sets, $\mathcal{V}$ is closed in $\mathbb{R}^{l}$.
@@
The set $\mathcal{V}\!\subset\!\mathbb{R}^{l}$ is not empty, i.e., system (1) features at least one steady-state admissible equilibrium point.
In essence, Assumption 1 establishes a continuous mapping from references to stabilizable equilibria, and Assumption 2 states that at least one of these equilibria is safe. By Assumption 1, let $\pi:\mathbb{R}^{n}\times\mathbb{R}^{l}\to\mathbb{R}^{m}$, locally Lipschitz continuous on $\mathbb{R}^{n}$ and continuous on $\mathbb{R}^{l}$, be a prestabilizing control law such that $\pi\big{(}\bar{x}(\boldsymbol{\mathbf{v}}),\boldsymbol{\mathbf{v}}\big{)}=\bar{u}(\boldsymbol{\mathbf{v}})$, and $\bar{x}(\boldsymbol{\mathbf{v}})$ is an asymptotically stable equilibrium point of the prestabilized system
$\dot{\boldsymbol{\mathbf{x}}}=f_{\pi}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\triangleq f(\boldsymbol{\mathbf{x}})+g(\boldsymbol{\mathbf{x}})\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}).$ (17)
Given a reference $\boldsymbol{\mathbf{v}}\in\mathbb{R}^{l}$, the prestabilized dynamics $f_{\pi}$ ensure the existence of an open set $\mathcal{D}_{\boldsymbol{\mathbf{v}}}\subset\mathbb{R}^{n}$ that satisfies $\bar{x}(\boldsymbol{\mathbf{v}})\in\mathcal{D}_{\boldsymbol{\mathbf{v}}}$ and is such that
@@
The set $\mathcal{D}_{\boldsymbol{\mathbf{v}}}$ is a region of attraction for $\bar{x}(\boldsymbol{\mathbf{v}})$. Note that, for each reference $\boldsymbol{\mathbf{v}}\in\mathbb{R}^{l}$, the input constraint set $\mathcal{U}$ induces state constraints on the prestabilized system. This allows us to define the reference-dependent state constraint set
By continuity of $\pi$ on $\mathbb{R}^{l}$, it follows that $\mathcal{X}_{\boldsymbol{\mathbf{v}}}$ is closed in $\mathbb{R}^{n}$.
Let us now define the state-reference space $\mathbb{R}^{n}\times\mathbb{R}^{l}$ obtained by concatenating the state and reference vectors $[\boldsymbol{\mathbf{x}};~{}\boldsymbol{\mathbf{v}}]$.
@@
The state-reference region of attraction set
$\tilde{\mathcal{D}}\triangleq\{(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\mathbb{R}^{n}\times\mathbb{R}^{l}\mid\boldsymbol{\mathbf{x}}\in\mathcal{D}_{\boldsymbol{\mathbf{v}}}\},$ (20)
is open in $\mathbb{R}^{n}\times\mathbb{R}^{l}$.
We also define the state-reference constraint set
$\tilde{\mathcal{X}}\triangleq\{(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\mathbb{R}^{n}\times\mathbb{R}^{l}\mid\boldsymbol{\mathbf{x}}\in\mathcal{X}_{\boldsymbol{\mathbf{v}}},\boldsymbol{\mathbf{v}}\in\mathcal{V}\}.$ (21)
The following DSM definition satisfies the one proposed in [23, Definition 1], although it makes stronger assumptions by requiring a) differentiability rather than continuity and b) invariance rather than strong returnability.
@@
Let $\Delta:\tilde{\mathcal{D}}\to\mathbb{R}^{n_{c}}$ be continuously differentiable and define the set $\tilde{\mathcal{C}}=\{(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{D}}\mid\Delta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq 0\}$. The function $\Delta$ is called a dynamic safety margin (DSM) for $\pi$ if the following conditions hold
$\tilde{\mathcal{C}}\subset\tilde{\mathcal{X}},$ (22a)
[…]
$\Delta_{i}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})=0~{}\implies~{}\frac{\partial\Delta_{i}}{\partial\boldsymbol{\mathbf{x}}}f_{\pi}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq 0,$ (22c)
where the last implication holds for all $i\in\{1,\ldots,n_{c}\}$.
@@
$\begin{bmatrix}\dot{\boldsymbol{\mathbf{x}}}\\ \dot{\boldsymbol{\mathbf{v}}}\end{bmatrix}=\begin{bmatrix}f(\boldsymbol{\mathbf{x}})+g(\boldsymbol{\mathbf{x}})\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\\ \min_{i\in\{1,\ldots,n_{c}\}}\Delta_{i}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\rho(\boldsymbol{\mathbf{v}},\boldsymbol{\mathbf{r}})\end{bmatrix},$ (23)
where $\rho:\mathbb{R}^{l}\times\mathbb{R}^{l}\to\mathbb{R}^{l}$ is any locally Lipschitz continuous function such that $\boldsymbol{\mathbf{r}}\in\mathbb{R}^{l}$ is an asymptotically stable equilibrium point of the system $\dot{\boldsymbol{\mathbf{v}}}(t)=\rho\big{(}\boldsymbol{\mathbf{v}}(t),\boldsymbol{\mathbf{r}}\big{)}$. In the ERG literature, $\rho$ is called a navigation field [23, Definition 2]. System (23) ensures that, if $\big{(}\boldsymbol{\mathbf{x}}(0),\boldsymbol{\mathbf{v}}(0)\big{)}\in\tilde{\mathcal{C}}$, then $\big{(}\boldsymbol{\mathbf{x}}(t),\boldsymbol{\mathbf{v}}(t)\big{)}\in\tilde{\mathcal{C}}\subset\tilde{\mathcal{X}}$ at all times [23]. The ERG literature provides systematic tools for computing DSMs for various classes of systems [30, 23, 31, 32, 33, 34]. The following section shows how these tools from the ERG literature can be used to construct valid CBFs.